# Beyond Chinchilla-Optimal: Accounting for Inference in Language Model Scaling Laws

### 31 Dec 2023 v1

@@ Abstract
Large language model (LLM) scaling laws are empirical formulas that estimate changes in model quality as a result of increasing parameter count and training data. However, these formulas, including the popular DeepMind Chinchilla scaling laws, neglect to include the cost of inference. We modify the Chinchilla scaling laws to calculate the optimal LLM parameter count and pre-training data size to train and deploy a model of a given quality and inference demand. We conduct our analysis both in terms of a compute budget and real-world costs and find that LLM researchers expecting reasonably large inference demand (~1B requests) should train models smaller and longer than Chinchilla-optimal.
@@ 1 Introduction
Large language models (LLM) have substantial training and inference compute and energy costs Knight (2023); Pope et al. (2022). Training computation costs are primarily determined by the size of the model and the amount of data it sees during training Hoffmann et al. (2022). State-of-the-art models have tens of billions of parameters and are trained on trillions of tokens Touvron et al. (2023b). Inference costs depend on the size of the model and the volume of user queries over the lifetime of the model. This volume can be significant: Demand for popular models can exceed billions of tokens per day OpenAI and Pilipiszyn (2021); Shazeer and Freitas (2022).
@@
Significant prior research has been conducted to find scaling laws, empirical formulas that estimate how changes in model and training data size impact model quality (Kaplan et al., 2020; Hoffmann et al., 2022). Hoffmann et al. (2022) is perhaps the most influential of these works; finding that to scale language models most efficiently, parameters and tokens should grow approximately equally. The authors applied this scaling law to train a 70B parameter model, Chinchilla, that outperformed much larger, more expensive models, including GPT-3. Subsequent LLMs have been trained following the Chinchilla scaling laws (Dey et al., 2023; Muennighoff et al., 2023).
@@
However, the Chinchilla scaling laws only account for the computational costs of training. By contrast, the LLaMA and LLaMA-2 family of models were trained on 1-2 trillion tokens, far more data than the Chinchilla scaling laws would deem “optimal” (Touvron et al., 2023a, b). Since inference costs are lower for smaller models, the extra training compute required to train a LLaMA-style model over a Chinchilla-style model of equivalent quality pays off after enough inference requests.
@@
Prior work has discussed the training-inference compute trade-off (Touvron et al., 2023a, b; Tow et al., 2023; De Vries, 2023; Villalobos and Atkinson, 2023). Touvron et al. (2023a) cites the lower inference cost of smaller models as inspiration for the LLaMA series. De Vries (2023) calculates the compute overhead of training longer than Chinchilla, but does not discuss quantify compute savings from inference. Recently, Villalobos and Atkinson (2023) discusses this trade-off in more detail, but shows the shift in scaling laws for only a single particular number of inferences.
@@
In this paper, we modify Chinchilla scaling laws to account for inference costs, calculating the optimal parameter and training token counts—both in terms of compute and dollar costs—to train and deploy a model of any given quality and inference demand. Our principled derivation estimates that LLM practitioners expecting significant demand (~$10^{9}$ inference requests) should train models substantially smaller and longer than Chinchilla-optimal.
@@ 2 Computational Optimality
The Chinchilla paper derived the parametric loss function in Eq. 1 and fit values for $A,B,E,\alpha$, and $\beta$ from the authors’ empirical training results. The best-fit values for these constants depend on the exact dataset and model architecture; however, the Chinchilla paper found largely consistent results across the MassiveText, Github (Rae et al., 2022), and C4 (Raffel et al., 2023) datasets, and subsequent work has replicated these scaling laws on other internet corpora and transformer variants Dey et al. (2023). Thus, we use the constant values from the Chinchilla paper in our analysis.
@@
$N^{*}(\ell,D_{\text{inf}}),D_{\text{tr}}^{*}(\ell,D_{\text{inf}})=\operatorname*{arg\,min}_{N,D_{\text{tr}}\mid L(N,D_{\text{tr}})=\ell}\operatorname*{T_{\text{FLOPs}}}(N,D_{\text{tr}})+\sum_{i}\operatorname*{I_{\text{FLOPs}}}(N,D_{\text{inf}}^{(i)}).$ (2)
@@
$N^{*}(\ell,D_{\text{inf}}),D_{\text{tr}}^{*}(\ell,D_{\text{inf}})=\operatorname*{arg\,min}_{N,D_{\text{tr}}\mid L(N,D_{\text{tr}})=\ell}6ND_{\text{tr}}+2ND_{\text{inf}}.$ (3)
@@
In Figure 1, we show how our inference-adjusted model’s FLOP counts, parameters, and pre-training tokens compare to Chinchilla-style models across a range of loss values and inference demands. When inference usage is significantly less than the number of pre-training tokens, Chinchilla models are essentially compute-optimal. However, as demand increases, inference costs becomes a significant factor. For a Chinchilla-7B-quality model with an inference demand of $10^{11}$ tokens, our formula suggests the compute-optimal method is to train a 6B parameter model on 1.18$\times$ the original data. For higher quality (i.e. larger and longer) models, the volume of inference demand required to shift the scaling law increases: An LLM developer that expects a 30B-Chinchilla-quality model will see $10^{13}$ tokens during inference can reduce their total FLOPs by 28% by training a 13.6B model on 2.84$\times$ the data. We provide additional results in Sec. B.1 in the Appendix.
@@ 3 Estimating Real-World Cost Optimality
Optimizing purely for minimum FLOPs has significant drawbacks which limit the applicability of our analysis in Section 2 to real-world deployments. The real-world cost of an inference request of $3D$ tokens is generally different than the cost to train on $D$ tokens. For instance, inference hardware utilization can be much lower than training utilization, since small batch size computation can result in low Model FLOPs Utilization (MFU). MFU can be as low as ~1% for inference (Pope et al., 2022) but is typically 40-60% during training (Korthikanti et al., 2022). Utilization is also different for input tokens vs. output tokens — since input tokens (prompts) are typically processed in a single forward pass, utilization is typically near training levels. By contrast, during generation, output tokens must be produced sequentially, resulting in low utilization due to memory bandwidth constraints. Another complicating factor is that inference operations can sometimes be cheaper than training FLOPs, since models can be quantized before inference time, turning 16- or 32-bit floating-point operations into 4- or 8-bit integer operations which run more efficiently on the same hardware. Quantization can also enable LLMs to fit on GPUs with less VRAM, so training and inference may occur on different hardware altogether (Frantar et al., 2023).
Figure 2: Ratios of (a) total cost, (b) model parameters, and (c) pre-training tokens, for cost-optimal models via our real-world estimation method vs. Chinchilla-style models. Results in this figure are shown with the following settings: training with 50% MFU, inference input with 50% MFU, generation with 1% MFU. Inference requests have 70 input tokens and 215 output tokens each, aligning with averages from real-world data (Zheng et al., 2023). To mimic a realistic scenario, we calculate costs assuming training occurs on A100-80GB and inference occurs on A100-40GB accelerators after INT8 quantization (see Sec. B.3 for details).
@@
$N^{*}(\ell,D_{\text{inp}},D_{\text{out}})^{*}(\ell,D_{\text{inp}},D_{\text{out}})$ $=\operatorname*{arg\,min}_{N,D\mid L(N,D_{\text{tr}})=\ell}\Bigg{[}\frac{C_{\text{tr}}}{\operatorname*{MFU_{\text{tr}}}}\operatorname*{T_{\text{FLOPs}}}(N,D_{\text{tr}})$ (4)
$+\sum_{i}\frac{C_{\text{inf}}}{\operatorname*{MFU_{\text{inp}}}}\operatorname*{I_{\text{FLOPs}}}(N,D_{\text{inp}}^{(i)})+\sum_{i}\frac{C_{\text{inf}}}{\operatorname*{MFU_{\text{out}}}}\operatorname*{I_{\text{FLOPs}}}(N,D_{\text{out}}^{(i)})\Bigg{]}.$ (5)
@@
$N^{*}(\ell,D_{\text{inp}},D_{\text{out}}),D_{\text{tr}}^{*}(\ell,D_{\text{inp}},D_{\text{out}})$ $=\operatorname*{arg\,min}_{N,D_{\text{tr}}\mid L(N,D_{\text{tr}})=\ell}\frac{6ND_{\text{tr}}C_{\text{tr}}}{\operatorname*{MFU_{\text{tr}}}}+2NC_{\text{inf}}\bigg{[}\frac{D_{\text{inp}}}{\operatorname*{MFU_{\text{inp}}}}+\frac{D_{\text{out}}}{\operatorname*{MFU_{\text{out}}}}\bigg{]}$ (6)

### 18 Jul 2024 v2

@@ Abstract
Large language model (LLM) scaling laws are empirical formulas that estimate changes in model quality as a result of increasing parameter count and training data. However, these formulas, including the popular Deepmind Chinchilla scaling laws, neglect to include the cost of inference. We modify the Chinchilla scaling laws to calculate the optimal LLM parameter count and pre-training data size to train and deploy a model of a given quality and inference demand. We conduct our analysis both in terms of a compute budget and real-world costs and find that LLM researchers expecting reasonably large inference demand (~1B requests) should train models smaller and longer than Chinchilla-optimal. Furthermore, we train 47 models of varying sizes and parameter counts to validate our formula and find that model quality continues to improve as we scale tokens per parameter to extreme ranges (up to 10,000). Finally, we ablate the procedure used to fit the Chinchilla scaling law coefficients and find that developing scaling laws only from data collected at typical token/parameter ratios overestimates the impact of additional tokens at these extreme ranges.
Scaling Laws, LLMs, Language Models, Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have substantial training and inference compute and energy costs (Knight, 2023; Pope et al., 2022). Training computation costs are primarily determined by the size of the model and the amount of data seen during training (Hoffmann et al., 2022). For state-of-the-art models with tens of billions of parameters trained on trillions of tokens, training costs can easily exceed millions of dollars. Similarly, inference costs depend on the size of the model and the volume of user queries over the lifetime of the model. This volume can be significant; demand for popular models can exceed billions of tokens per day (OpenAI & Pilipiszyn, 2021; Shazeer & Freitas, 2022).
@@
Recent studies have proposed scaling laws, empirical formulas that estimate how changes in model and training data size impact model quality (Kaplan et al., 2020; Hoffmann et al., 2022). Hoffmann et al. (2022) is perhaps the most influential of these works, finding that to scale language models most efficiently, parameters and tokens should grow approximately linearly. The authors applied this scaling law to train a 70B parameter model (dubbed Chinchilla) that outperformed much larger and more expensive models such as GPT-3. As a result, many subsequent LLMs have been trained following the Chinchilla scaling laws (Dey et al., 2023; Muennighoff et al., 2023).
@@
However, the Chinchilla scaling laws only account for the computational costs of training. By contrast, the Llama 2 family of models were trained on 2 trillion tokens and the Llama 3 family of models were trained on 15 trillion tokens, which is far more data than the Chinchilla scaling laws would deem “optimal” (Touvron et al., 2023a, b; AI@Meta, 2024). Since inference costs are lower for smaller models, the extra training compute required to train a Llama-style model over a Chinchilla-style model of equivalent quality pays off after enough inference requests.
Figure 1: Schematic of compute savings achieved via our approach. An LLM developer seeking to train a 13B model who expects 2 trillion tokens of inference demand during the model’s lifetime can reduce their total compute by $1.7\times 10^{22}$ FLOPs ($17\%$) by instead training a 7B model on more data (a). The extra compute required to train the 7B model beyond its Chinchilla-optimal point to match the 13B’s quality is made up for during inference (b), (c). Our method quantifies this training-inference trade-off, producing models that are optimal over their total lifetime.
@@
Prior work has discussed the training-inference compute trade-off (Touvron et al., 2023a, b; Tow et al., 2023; De Vries, 2023; Villalobos & Atkinson, 2023). Touvron et al. (2023a) cites the lower inference cost of smaller models as inspiration for the LLaMA series. De Vries (2023) calculates the compute overhead of training longer than Chinchilla, but does not discuss quantify compute savings from inference. Villalobos & Atkinson (2023) discuss this trade-off in more detail, but show the shift in scaling laws for only a single particular number of inferences.
Other related work includes Muennighoff et al. (2023), which adapts the Chinchilla scaling laws for the data-constrained regime, where we have to repeat training tokens. Our problem setting is the opposite: We assume we are data rich but compute constrained, and seek to minimize computation costs assuming we have enough data to train high-quality models.
@@
In this paper, we modify the Chinchilla scaling laws to account for inference costs by calculating the optimal parameter and training token counts—both in terms of compute (Sec. 2) and dollar costs (Sec. 6)—to train and deploy a model of any given quality and inference demand. Our principled derivation estimates that LLM practitioners expecting significant demand (~$10^{9}$ inference requests) should train models substantially smaller and longer than Chinchilla-optimal. Figure 1 illustrates the benefits of our compute-optimal method for a realistic scenario.
In inference-heavy regimes, our modification predicts that it is optimal to train a model far smaller and longer than Chinchilla predicts, up to thousands of tokens per parameter. Does this hold true in practice? Do transformer models see continued improvements at such extreme cases? Or is there a point beyond which models saturate, and additional tokens provide no further improvement, as De Vries (2023) suggests? To uncover the behavior of transformer models in these cases, we train 47 models ranging from 150M to 6B parameters, on various data budgets from 10 to 10,000 tokens per parameter. We find that model quality continues to improve as we scale token ratios. We do not find evidence of a “saturation point,” beyond which models do not improve, even with additional data.
Lastly, we ablate the parametric curve fitting procedure from the Chinchilla scaling laws. The Chinchilla scaling laws use empirical data from over 400 training runs to determine coefficients that estimate precisely how additional model parameters and training data impact loss. Hoffmann et al. (2022) collected data only from standard token ratio training runs ($\leq$~100 tokens/parameters). Our ablation indicates that when fitting Chinchilla coefficients using only typical token ratio runs, this formula overestimates the impact of additional training data as we move to the extreme ratio regime.
@@ 2 Computational Optimality
The Chinchilla paper derived the parametric loss function in Eq. 1 and fit values for $A,B,E,\alpha$, and $\beta$ from the authors’ empirical training results. The best-fit values for these coefficients depend on the exact dataset and model architecture; however, the Chinchilla paper found largely consistent results across the MassiveText, Github (Rae et al., 2022), and C4 (Raffel et al., 2023) datasets, and subsequent work has replicated these scaling laws on other internet corpora and transformer variants (Dey et al., 2023; Besiroglu et al., 2024; Gadre et al., 2024). We use the coefficient values from the Chinchilla paper in our analysis here, and explore fitting these coefficients on different datasets and data ratios more detail in Section 5.
@@
$\begin{split}N^{*}(\ell,D_{\text{inf}}),D_{\text{tr}}^{*}(\ell,D_{\text{inf}})=&\operatorname*{arg\,min}_{N,D_{\text{tr}}\mid L(N,D_{\text{tr}})=\ell}\operatorname*{T_{\text{FLOPs}}}(N,D_{\text{tr}})\\ &+\sum_{i}\operatorname*{I_{\text{FLOPs}}}(N,D_{\text{inf}}^{(i)}).\end{split}$ (2)
@@
$\begin{split}N^{*}(\ell,D_{\text{inf}}),D_{\text{tr}}^{*}(\ell,D_{\text{inf}})&=\\ &\operatorname*{arg\,min}_{N,D_{\text{tr}}\mid L(N,D_{\text{tr}})=\ell}6ND_{\text{tr}}+2ND_{\text{inf}}.\end{split}$ (3)
@@
Figure 3: For each model (150M, 370M, 750M, 1.3B, 2.5B, 6B) in our experimental sweep, we plot (a) Loss vs. Tokens per parameter, (b) Gauntlet Average (an aggregation of all our metrics described in Sec. 3) vs. Tokens per parameter, and (c) Loss vs. Gauntlet Average. Category-specific Gauntlet results are available in Sec. D.
@@
In Figure 2, we show how our inference-adjusted model’s FLOP counts, parameters, and pre-training tokens compare to Chinchilla-style models across a range of loss values and inference demands. When inference usage is significantly less than the number of pre-training tokens, Chinchilla models are essentially compute-optimal. However, as demand increases, inference costs becomes a significant factor. For a 7B-Chinchilla-quality model with an inference demand of $10^{11}$ tokens, our formula suggests the compute-optimal method is to train a 6B parameter model on 1.18$\times$ the original (Chinchilla-prescribed) amount of data. For higher quality models (i.e. models that are larger and/or trained for longer), the volume of inference demand required to shift the scaling law increases: An LLM developer that expects a 30B-Chinchilla-quality model will see $10^{13}$ tokens during inference can reduce their total FLOPs by 28% by training a 13.6B model on 2.84$\times$ the data. We provide additional results in Sec. B.1 in the Appendix.
3 Experiments
In high-demand inference scenarios, our analysis in Section 2 suggests that we should train models significantly smaller and on much more data than Chinchilla-style, resulting in hundreds or even thousands of tokens per parameter. The Chinchilla scaling laws have no upper bound—their form infinitely predicts that as tokens per parameter increase, model quality increases. However, the Chinchilla authors do not validate the scaling law at the outer ranges, conducting experiments only at typical ($<$~100 tokens/parameter) ratios. De Vries (2023) postulates that there is a critical model size (~30%), below which, it is not possible to train on any number of tokens and match a Chinchilla-style model.
To characterize the behavior of transformers at extreme data sizes, we train 47 models with the MPT architecture (MosaicML, 2023) of varying size and token ratios. Our models range from 150M to 6B parameters, and our data budgets from 10 to 10,000 tokens per parameter. Due to resource constraints, we could not complete a full sweep for all model sizes (e.g. we train our 2.5B model up to 500 tokens/parameter). Our dataset consists of trillions of tokens of general web text and code. For all experiments, we train for only a single epoch and do not repeat data. Further details are provided in Section C in the Appendix.
Furthermore, to ensure that loss is a good proxy for downstream metrics, we evaluate each model on a version of the open source Evaluation Gauntlet (MosaicML NLP Team, 2023), with tasks in five categories:
  • World Knowledge: This includes Jeopardy (Wolfe et al., 2022), MMLU (Hendrycks et al., 2020), BIG-bench WikiData (Srivastava et al., 2022), ARC Easy and ARC Challenge (Clark et al., 2018).
  • Commonsense Reasoning: BIG-bench Strategy QA and BIG-bench Strange Stories (Srivastava et al., 2022), COPA (Roemmele et al., 2011), PIQA (Bisk et al., 2020), OpenBook QA (Mihaylov et al., 2018), Common Sense QA (Talmor et al., 2018), and SIQA (Sap et al., 2019).
  • Reading Comprehension: SQuAD (Rajpurkar et al., 2016), BoolQ (Clark et al., 2019), CoQA (Reddy et al., 2019), and AGI Eval (Zhong et al., 2023).
  • Language Understanding: LAMBADA (Paperno et al., 2016), HellaSwag (Zellers et al., 2019), Winograd Schema Challenge (Levesque et al., 2012), and Winogrande (Sakaguchi et al., 2019).
  • Symbolic Problem Solving: BIG-bench Elementary Math QA (Srivastava et al., 2022) BIG-bench Dyck Languages (Srivastava et al., 2022), BIG-bench Operators (Srivastava et al., 2022), Math QA (Amini et al., 2019), LogiQA, GSM8k (Cobbe et al., 2021), SVAMP (Patel et al., 2021), AGI Eval SAT Math and AGI Eval LSAT (Zhong et al., 2023).
To compute model performance on the above datasets, the Evaluation Gauntlet uses the InContextLearningQAAccuracy for question-answering tasks, InContextLearningLMAccuracy for language-modeling tasks, and InContextLearningMultipleChoiceAccuracy for multiple-choice tasks. All metrics are from the Composer library (The Mosaic ML Team, 2021).
The Gauntlet Average is an average of the accuracy of all the above tasks, with each task weighted equally after subtracting out the task’s baseline (random) accuracy and normalizing.
4 Results
In Figure 3, we show the major results of our experiments. Our key finding is that loss continues to decrease as we increase tokens per parameter, even to extreme ratios. Although it takes exponentially more tokens to reduce loss at large ratios, loss does not plateau as we scale to 10,000 tokens per parameter for our 150M model. For models larger than 150M, we only test up to a maximum of 1,000 tokens per parameter due to resource constraints, and at this scale we also see no evidence of loss flat-lining. Further experiments are needed to see how loss scales beyond this point.
Existing literature claims there exists a critical model size, below which we cannot match a Chinchilla-optimal model’s quality (De Vries, 2023). Since our experiments do not see any plateauing of loss as we scale token ratios, we find no evidence to support the critical size hypothesis, although further testing is needed at extreme scales—it is possible that beyond 10,000 tokens/parameter, behavior changes.
We also show results for downstream metrics. Fig. 3(b) shows that the Gauntlet Average improves as we increase tokens per parameter — again, we find no evidence of a “saturation” point, beyond which additional tokens do not result in better performance. In fact, as Fig. 3(c) shows, as loss decreases, smaller decreases in loss lead to larger improvements in downstream accuracy.
In Fig. 3(c), we plot the Gauntlet Average as a function of loss. Loss and Gauntlet Average are tightly correlated, showing that improvements in loss are excellent predictors for improvements in general model quality. Our results show that practitioners interested in predicting downstream metrics as a function of model parameters and token counts can make use of existing scaling laws to accurately understand how their downstream metrics change at scale.
In Figure 4, we plot loss vs. FLOPs, grouping our data points by token-per-parameter ratio (up to 500 tokens/param) instead of model size, following Gadre et al. (2024). For token-per-parameter ratios $\geq 20$, our lines of best fit are nearly parallel, indicating that models learn with similar efficiency. This result is striking, as it indicates that loss decreases at similar rates for each additional FLOP regardless of whether training occurs in the standard Chinchilla regime or at extreme ratios (500 tokens/param). Our results also confirm that training at low ($<20$) token-per-parameter ratios is less computationally efficient, in line with the results from the Chinchilla paper.
In Section D in the Appendix, we show results for each Gauntlet category. Since they consist of fewer tasks, category averages show less consistent correlation with data and model sizes than the overall Gauntlet Average, because on many tasks smaller models are not able to achieve significantly better performance than a random baseline (Barton, 2024).
Figure 4: Log-log plot of loss vs. FLOPs, grouped by token/parameter ratio. The slope of each trendline shows how efficiently models learn at each ratio.
5 Parametric Fitting
Figure 5: Chinchilla parameteric scaling curves fit to progressively larger subsets of our empirical training data. The coefficients that define these curves are detailed in Table 1.
The Chinchilla scaling law has five coefficients ($A,B,E,\alpha$ and $\beta$) which were found empirically by fitting Eq. 1 to data collected from 400 training runs by Hoffmann et al. (2022). These coefficients determine precisely how much the scaling law weighs increasing model parameters vs. increasing training data. Muennighoff et al. (2023) found similar coefficient values after training 54 similarly-sized runs on C4, one of the datasets tested in the Chinchilla paper. In both cases, the empirical data collected was largely from small token-to-parameter training runs ($<$~100). Since Hoffmann et al. (2022) and Muennighoff et al. (2023) focused on optimizing training compute and data efficiency, respectively, it is reasonable that they concentrated empirical data collection around the training efficiency “sweet spot,” around 20 tokens per parameter. However, when we incorporate inference costs into the picture, we may need to train models for hundreds or thousands of tokens per parameter. Thus, we require scaling laws that can accurately predict training loss far beyond typical data ratios.
To understand how well the Chinchilla parametric fitting procedure generalizes to loss values for extremely long training runs, we fit parametric curves on successively larger subsets of our training data. First, we use only training runs with no more than 100 tokens per parameter, followed by runs with no more than 250 tokens per parameter, etc. For each of the curves, we follow the procedure described in the Sec. D.2 of the Chinchilla paper, minimizing the objective in Eq. 4, where LSE is the log-sum-exp operator, and $\delta=10^{-3}$. We use the L-BFGS algorithm to perform the minimization, initialized from a grid of starting points for each of the parameters. We then set $A,B,E=\exp(a),\exp(b),\exp(e)$ and minimize the Huber loss in the following equation:
$\begin{split}\min_{a,b,e,\alpha,\beta}\sum_{\text{Run }i}\text{Huber}_{\delta}\Big{(}\text{LSE}\big{[}&a-\alpha\log(N^{(i)}),\\ &b-\beta\log(D_{\text{tr}}^{(i)}),e\big{]}-\log(L^{(i)})\Big{)}\end{split}$ (4)
Our results are shown in Table 1 and visualized in Figure 5. First, we see that our parameteric curves follow a consistent trend—as we use more extreme data in our fitting procedure, our scaling curves become flatter. This trend suggests that if we only use data from typical token ratios to determine our scaling law coefficients, we will overestimate the impact of additional training data as we move towards the long-data-ratio regime. Taken together with our results from Sec. 4, our experiments show that although models continue to learn at extreme training durations, they do so more slowly than scaling laws predict.
Second, somewhat surprisingly, none of of our parametric curves fit our 150M long-ratio training results well. It appears that as we extend training duration far beyond typical Chinchilla ratios, the parametric loss function is not flexible enough to accurately model the behavior of both smaller ($\leq$150M) and larger models. Further research is needed to understand the limits of scaling law extrapolation.
We note that our experimentally-derived coefficients differ from those found in the Chinchilla paper. This is expected, as both the data and the model architectures we used to train all checkpoints are quite different from the original work (Hoffmann et al., 2022). Besiroglu et al. (2024) noted in a replication study that the coefficients reported in Hoffmann et al. (2022) were rounded in ways that leads to significant bias in scaling law predictions.
Table 1: Parametric fitting on data including extreme training runs. The Chinchilla coefficients from Hoffmann et al. (2022) are included for reference.
Data $\alpha$ $\beta$ A B E
$\leq$100 tok/param 0.08 0.13 7.199 25.97 0.17
$\leq$250 tok/param 0.13 0.16 14.23 39.54 0.98
$\leq$500 tok/param 0.13 0.16 17.07 35.80 0.95
All Data 0.18 0.24 33.66 138.9 1.45
Chinchilla 0.34 0.28 406.4 410.7 1.69
Since Hoffmann et al. (2022) trained almost exclusively on small-duration training runs ($\leq$~100 tokens/parameter), our results suggest that that in their original form, the Chinchilla scaling laws do not extend to extreme-duration training runs. To the extent that they do, current scaling laws overestimate the improvements in loss that stem from long-duration training on additional data.
These results have significant impact as the field continues to train longer and longer. For example, LLaMA 7B was trained on 1 trillion tokens (Touvron et al., 2023a), Llama 2 7B was trained on 2 trillion tokens (Touvron et al., 2023b), and Llama 3 8B was trained on 15 trillion tokens (AI@Meta, 2024). How much of the quality difference between these models is due to simply training on more data? How much is due to other changes, like architecture modifications or data quality improvements? Our results indicate that if we apply the Chinchilla scaling laws to understand the scaling component of this quality improvement, we will likely overestimate the its impact of more data versus the other changes.
Further work is needed to fully characterize scaling laws at extreme token to parameter ratios. Due to resource constraints, we do not collect data at the same scale as the Chinchilla paper—both in terms of model size (we only test up to 6B vs. 16B), and number of training runs (47 vs. 400). Little research has been conducted to explore the minimum data required to accurately fit scaling laws coefficients (see Besiroglu et al. (2024)).
Figure 6: Ratios of (a) total training + inference cost, (b) model parameters, and (c) pre-training tokens, for cost-optimal models via our real-world estimation method vs. Chinchilla-style models. Results in this figure are shown with the following settings: training with 50% MFU, inference input with 50% MFU, generation with 1% MFU. Inference requests have 70 input tokens and 215 output tokens each, aligning with averages from real-world data (Zheng et al., 2023). To mimic a realistic scenario, we calculate costs assuming training occurs on A100-80GB and inference occurs on A100-40GB accelerators after INT8 quantization (see Sec. B.3 for details).
@@ 6 Estimating Real-World Cost Optimality
Our method introduced in Section 2 optimizes purely for minimum total (training + inference) FLOPs. However, this has significant drawbacks which limit its applicability to real-world deployments. The real-world cost of an inference request of $3D$ tokens is generally different than the cost to train on $D$ tokens. For instance, inference hardware utilization can be much lower than training utilization, since small batch size computation can result in low Model FLOPs Utilization (MFU). MFU can be as low as ~1% for inference (Pope et al., 2022) but is typically 40-60% during training (Korthikanti et al., 2022). Utilization is also different for input tokens vs. output tokens — since input tokens (prompts) are typically processed in a single forward pass, utilization is typically near training levels. By contrast, during generation, output tokens must be produced sequentially, resulting in low utilization due to memory bandwidth constraints. Another complicating factor is that inference operations can sometimes be cheaper than training FLOPs, since models can be quantized before inference time, turning 16- or 32-bit floating-point operations into 4- or 8-bit integer operations which run more efficiently on the same hardware. Quantization can also enable LLMs to fit on GPUs with less VRAM, so training and inference may occur on different hardware altogether (Frantar et al., 2023).
@@
$\begin{split}N^{*}(\ell,D_{\text{inp}},D_{\text{out}}),&D_{\text{tr}}^{*}(\ell,D_{\text{inp}},D_{\text{out}})=\\ &\operatorname*{arg\,min}_{N,D\mid L(N,D_{\text{tr}})=\ell}\Bigg{[}\frac{C_{\text{tr}}}{\operatorname*{MFU_{\text{tr}}}}\operatorname*{T_{\text{FLOPs}}}(N,D_{\text{tr}})\\ &+\sum_{i}\frac{C_{\text{inf}}}{\operatorname*{MFU_{\text{inp}}}}\operatorname*{I_{\text{FLOPs}}}(N,D_{\text{inp}}^{(i)})\\ &+\sum_{i}\frac{C_{\text{inf}}}{\operatorname*{MFU_{\text{out}}}}\operatorname*{I_{\text{FLOPs}}}(N,D_{\text{out}}^{(i)})\Bigg{]}.\end{split}$ (5)
@@
$\begin{split}&N^{*}(\ell,D_{\text{inp}},D_{\text{out}}),D_{\text{tr}}^{*}(\ell,D_{\text{inp}},D_{\text{out}})=\\ &\operatorname*{arg\,min}_{N,D_{\text{tr}}\mid L(N,D_{\text{tr}})=\ell}\frac{6ND_{\text{tr}}C_{\text{tr}}}{\operatorname*{MFU_{\text{tr}}}}+2NC_{\text{inf}}\bigg{[}\frac{D_{\text{inp}}}{\operatorname*{MFU_{\text{inp}}}}+\frac{D_{\text{out}}}{\operatorname*{MFU_{\text{out}}}}\bigg{]}\end{split}$ (6)
@@
7 Related Work
Many studies have contributed to the development of scaling laws for LLMs, including Hestness et al. (2017, 2019), Rosenfeld et al. (2019), Henighan et al. (2020), Kaplan et al. (2020), Sorscher et al. (2022), Tay et al. (2022), and Caballero et al. (2022) (see Villalobos (2023) for a review). Some of these studies focused on scaling laws for transfer settings (i.e. downstream performance), such as Hernandez et al. (2021); Mikami et al. (2021); Abnar et al. (2021) and Tay et al. (2022).
A few studies have also (gently) critiqued the general parametric function fitting approach of Hoffmann et al. (2022). Besiroglu et al. (2024) attempted to replicate the methodology used in Hoffmann et al. (2022) and found that the confidence intervals reported in the original study were implausibly narrow. The broader implication is that confidence intervals are quite wide for parametric function fitting on a small number of data points. This is also a potentially valid critique of our empirical results, as our analysis only includes 47 separate experiments (Chinchilla included more than 400 experiments). In a similar vein, Porian et al. (2024) investigate discrepancies between the scaling laws from Kaplan et al. (2020) and Hoffmann et al. (2022).
A handful of compelling scaling law papers have been published since 2023, when an earlier version of this work was first presented (Sardana & Frankle, 2023). For example, Krajewski et al. (2024) characterize differences in scaling properties between dense transformers and Mixture of Expert (MoE) models. More theoretical studies include Michaud et al. (2024) and Bordelon et al. (2024). Paquette et al. (2024) uses phase-plane analysis to characterize scaling laws in the compute-limited, infinite-data regime. Ruan et al. (2024) builds scaling laws from multiple LLM model families using low-dimensionality tools applied to publicly available data of LLM performance on open benchmarks.
The results presented in Gadre et al. (2024) are particularly relevant to this paper. The authors train 100 models between the sizes of 1.4B and 6.9B parameters and on data with tokens-per-parameter ratios between 20 and 640. Similar to our study, they find reliable scaling laws in these model and data regimes. They also find that downstream task performance is strongly correlated to LLM perplexity.